A concentrated solution of HCl has a molarity of 7.75 M. You add 25.00 mL to a flask and then dilute with water. The mass of the final solution is 125.0 g. What is the mass percent of HCl ?
Answer #1

Dear student,

Answear Concentration HCl = 7.75M Volume = 25ml = 25x103 Lt moles of Hel = Molarity & volume = 7.75*25*10 3 = 0.194 mol = No.Now, % mass of Het = mass of HCl X luo massa final solution = 7. L x 100 125 - 0.0568 x 100 = 5.68).

Please give thumb's up to this answer.

With regards...